[A] - Pa(KH) H=0.033 TAJ = 0.033 12.0 atm) =10.066 M 3. Calculate the concentration of TIH(aq) in the three scenarios below: just B and C, Thanks!
Answer #1

b)

Balanced chemical equation:-

TlCl(s) <-----> Tl+(aq) + Cl-(aq)

It is evident from the balanced chemical equation that 1 mole of TlCl produces 1 mole of Tl+ and 1 mole of Cl-.

i.e., x moles of TlCl will produce x moles of Tl+ and x mole of Cl-.

Filling Ice table

TlCl Tl+ Cl-
I 0.012 0 0
C -x +x +x
E 0.012-x x x

Here,

I stands for Initial no. of moles.

C stands for change in no. of moles, -ve sign denotes loss and +ve sign denotes gain.

E stands for equilibrium/final no. of moles.

Finding concentration using the formula-

Concentration (M) = \tfrac{No. of moles}{Volume in litre} ------(a)

Volume of water = 1 L -----(Given)

Concentration of Tl+ = (x/1) ------using equation(a)

[Tl+] = x M

Concentration of Cl- = (x/1) ------using equation(a)

[Cl-] = x M

Expression for Ksp can be written as-

K_{sp} = \tfrac{[Tl^{+}][Cl^{-}]}{[TlCl]} ------(b)

Here,

Ksp = The solubility product constant

[Tl+] = Concentration of Tl+

[Cl-] = Concentration of Cl-

[TlCl] = Concentration of TlCl

In the given chemical reaction, TlCl is solid, the concentration of solid in expression for Ksp is taken as 1.

Therefore, [TlCl] = 1

Hence, equation(b) can be written as-

Ksp = [Tl+][Cl-] ------(b')

Ksp = 1.4X10^-4 ------(Given)

= 0.00014

Putting all values in equation(b'), we get,

0.00014 = x.x

0.00014 = x^2

x = \sqrt{0.00014}

x = 0.0118 M

[Tl+] = 0.0118 M

The concentration of Tl+ in the given scenario is 0.0118 M.

c)

Balanced chemical equation:-

TlCl(s) <-----> Tl+(aq) + Cl-(aq)

It is evident from the balanced chemical equation that 1 mole of TlCl produces 1 mole of Tl+ and 1 mole of Cl-.

i.e., x moles of TlCl will produce x moles of Tl+ and x mole of Cl-.

Filling Ice table

TlCl Tl+ Cl-
I 1.5 0 0
C -x +x +x
E 1.5-x x x

Here,

I stands for Initial no. of moles.

C stands for change in no. of moles, -ve sign denotes loss and +ve sign denotes gain.

E stands for equilibrium/final no. of moles.

Finding concentration using the formula-

Concentration (M) = \tfrac{No. of moles}{Volume in litre} ------(a)

Volume of water = 1 L -----(Given)

Concentration of Tl+ = (x/1) ------using equation(a)

[Tl+] = x M

Concentration of Cl- = (x/1) ------using equation(a)

[Cl-] = x M

Expression for Ksp can be written as-

K_{sp} = \tfrac{[Tl^{+}][Cl^{-}]}{[TlCl]} ------(b)

Here,

Ksp = The solubility product constant

[Tl+] = Concentration of Tl+

[Cl-] = Concentration of Cl-

[TlCl] = Concentration of TlCl

In the given chemical reaction, TlCl is solid, the concentration of solid in expression for Ksp is taken as 1.

Therefore, [TlCl] = 1

Hence, equation(b) can be written as-

Ksp = [Tl+][Cl-] ------(b')

Ksp = 1.4X10^-4 ------(Given)

= 0.00014

Putting all values in equation(b'), we get,

0.00014 = x.x

0.00014 = x^2

x = \sqrt{0.00014}

x = 0.0118 M

[Tl+] = 0.0118 M

The concentration of Tl+ in the given scenario is 0.0118 M.